Comment lampoly SNF 2010-11-04 16:57:58: Update

mtang at snf.stanford.edu mtang at snf.stanford.edu
Thu Nov 4 16:57:58 PDT 2010


Entrance loadlock door is stuck.  Cesar cleaned the pneumatics to get them working last time and had ordered replacement ones -- it's likely these will have to be replaced now.
Cesar adjusted the gross ICP tune position.  This morning, we ran about a dozen wafers on recipe 1 without problem although noticed that the load setting was creeping up to the maximum of 32525.  He adjusted the ICP tune position to 0 degrees and we ran more wafers.  After about a dozen, the RF tune problem occurred again, about 20 seconds into step 7 (same symptoms as yesterday, where system tunes nicely, but suddenly detunes after 75-90 seconds of plasma.)
Jim McV has some ideas and I hope to corner him tomorrow.




More information about the lampoly-pcs mailing list